OpenStudy (anonymous):

Let f(x)=3x^2+3x+4..what are these? 2f(a) f(2a) f(a^2) [f(a)]^2 f(a+h)

myininaya (myininaya):

\[f(a+h)=3(a+h)^2+3(a+h)+4 \] see how i just replaced all the x's in your expression with a+h

myininaya (myininaya):

\[f( * )=3*^2+3*+4\]

myininaya (myininaya):

\[f(tuna fish)=3(tuna fish)^2+3(tuna fish)+4\]

myininaya (myininaya):

that thing inside f( ) is what we are going to replace x with

OpenStudy (anonymous):

what about the first one, there is a 2 before the f. what do i do about it?

myininaya (myininaya):

\[2f(a)=2(f(a))=2(3a^2+3a+4)\]

myininaya (myininaya):

\[f(a^2)=3(a^2)^2+3(a^2)+4\]

myininaya (myininaya):

\[[f(a)]^2=[3a^2+3a+4]^2\]